OpenStudy (anonymous):

How do you find the slope of a line when given this equation... 15x-3y=90?

OpenStudy (anonymous):

y=mx+b

OpenStudy (amistre64):

ax +by = c slope = -a/b

OpenStudy (amistre64):

\(\frac{-(15)}{-3}\)

OpenStudy (anonymous):

Move the x term to the other side of the equation, divide the coefficient of the x term by the coefficient of the y term, and you have the slope.

OpenStudy (anonymous):

-3y=-15x+90 solve slope = 5